OpenStudy (anonymous):

Find the exact value of cot^-1(1).

OpenStudy (anonymous):

\[\frac{1}{\text{ArcCot}[1]}=\frac{4}{\pi } \]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[\text{Cot}[45{}^{\circ}]=1 \]\[45{}^{\circ}=\frac{\pi }{4} \]\[1\left/\left(\frac{\pi }{4}\right)\right.=\frac{4}{\pi } \]

OpenStudy (anonymous):

sorry for being clueless but how did you get cot(45)=1? I understand everything after that

OpenStudy (anonymous):

Check out the following site: http://www.mathwords.com/t/trig_values_of_special_angles.htm It is convenient to have access to a table of the special angles.

OpenStudy (anonymous):

Keep in mind that ArcCot[1] mean stands for, "the angle whose cotangent is equal to one."